OpenStudy (anonymous):

Let f(x) = x2 - 2x + 2 and g(x) = x - 3. Find f(x) • g(x). Question 5 options: 1) x3 - 6x2 + 12x - 6 2) x3 - 5x2 + 8x - 6 3) x3 - 3x2 + 6x - 6 4) x3 - 2x2 + 5x - 6

OpenStudy (lalaly):

(x^2-2x+2)(x-3) = x^3 -3x^2 -2x^2 +6x +2x -6 = x^3 -5x^2 +8x -6 which is 2)